Object A of charge -5e-9 C is located on the x-axis at x = 0.5 m. Point P is located on the x-axis at x = 0.6m. Object B of charge 1e-9 C is located on the x-axis at x = 1.1 m. Determine the magnitude and direction of the net electric field at point P due to objects A and B.

4989.6 N ⁄ C West
4536 N ⁄ C West
5443.2 N ⁄ C West
4536 N ⁄ C East
4989.6 N ⁄ C East
1
Expert's answer
2019-07-19T14:37:46-0400

The electric field at poit P


"E=E_1-E_2=k\frac{q_1}{AP^2}-k\frac{q_2}{BP^2}"

"=9\times 10^9\:\rm{N\cdot m^2/C^2}\left(\frac{-5\times 10^{-9}\:\rm{C}}{0.1^2\:\rm{m^2}}-\frac{1\times 10^{-9}\:\rm{C}}{0.5^2\:\rm{m^2}}\right)"

"=4536\:\rm{N/C}\quad\rm{East}"